OSCILLATOR STRENGTHS IN THE VISIBLE ABSORPTION SPECTRUM OF I_{2}

J. TELLINGHUISEN, Department of Chemistry, Vanderbilt University, Nashville, TN 37235.

The B − X transition in I_{2} stands as one of the most precisely characterized diatomic electronic systems, with thousands of lines having been measured with precision sufficient to serve as frequency standards across the 500-700-nm spectral region. Accordingly the molecular constants of the B − X system permit recalculation of those line positions with similar precision ( 0.001 cm^{−1}). Yet the B − X electronic transition strength remains a 10-percent quantity, in part from the difficulty of dealing with two overlapping continuous 1u-X transitions, A−X and C −X. In this work, oscillator strengths are estimated from a long-known but little used method – integration of absorption cross sections over single rotational lines – using measurements obtained for lines near 650 nm with a diode laser. The results are combined with new, precise low-resolution (1 nm) absorption data to obtain a refined assessment of electronic transition strengths in the I_{2} visible absorption spectrum.